OpenStudy (anonymous):

evaluate the definite integral of (x sin(x))/cos^3(x) dx, on the interval [0, pi/4]

zepdrix (zepdrix):

Hmm looks like `Integration by Parts` for this problem. \[\Large\rm \int\limits \frac{x \sin x}{\cos^3x}~dx\] \[\Large\rm u=x,\qquad\qquad dv=\frac{\sin x}{\cos^3x}~dx\]Understand how to find your v?

OpenStudy (anonymous):

eek - not from sinx/cos^3x dx :(

zepdrix (zepdrix):

u-sub if you need to. u=cos x You end up with an integral like 1/u^3, leading to 1/(2cos^2x) or something similiar.

zepdrix (zepdrix):

Err actually, I don't want to use u for our parts AND here. Let's call it something else, \[\Large\rm m=\cos x, \qquad\qquad -dm=\sin x~dx\] \[\Large\rm \int\limits \frac{-dm}{m^3}=\frac{1}{2m^2}=\frac{1}{2cos^2x}\]Like uhhh that? Yah?

OpenStudy (anonymous):

OK! Now, the integration by parts formula that I know is \[\int\limits u dv=uv - \int\limits vdu\] Is this the formula to use even though I have \[\frac{ u }{ dv }\]

OpenStudy (freckles):

You have u dv not u/dv Recall he called x=u and the other part of the fraction dv

OpenStudy (anonymous):

Oh - got it!

OpenStudy (anonymous):

Thank you so much!! I think I have an answer. :)